BIOACTIVITY GUIDED ISOLATION OF A NOVEL ANTI-INFLAMMATORY AND ANTIBACTERIAL FLAVONOL AND OXYCHROMONE FROM THE PLANT PILIOSTIGMA RETICULATUM (SCHUM)

\underline{DARAMOLA O. ABIMBOLA} and BABAJIDE O. OMOTOLA, Department of Chemistry, Obafemi Awolowo University, Adeyemi Campus, P.M.B 520, Ondo, Ondo State Nigeria; BABAJIDE J. OKALEKAN, Department of Chemistry,University of the Western Cape, Private Bag X17, Bellville, 7535, South Africa.

Piliostigma reticulatum (Schum) is a plant used throughout tropical Africa in treating varieties of ailments, which includes ulcers, cough, wounds, chest pains, gingivitis, fever, gonorrhea etc.In a screening of fifteen plant species from an inventory of some African medicinal plants for anti-inflammatory activity, Piliostigma reticulatum leaf extract was found to possess highly moderate but specific prostaglandin synthesis inhibitory activity. The plant extract was also found to exhibit antibacterial activity against Staphylococcus aureus NCTC 10788, Escherichia coli NCTC 9001, Bacillus subtilis NCTC 8236 and Proteus vulgaris NCTC 4175. This observation prompts in looking into the constituents present in Piliostigma reticulatum leaves with its widespread ethnomedicinal uses. A total of Ten compounds were isolated and their structures were unambiguously established by spectroscopic methods including infrared and ultraviolet spectroscopy, high-resolution mass spectrometry and Nuclear magnetic resonance spectrometry. Four of the isolated compounds were Novels. This includes the first ever P1 C-Methyl-p-phenoxychromonol (Piliostigmol), P2 6,8-Di-C-Methylquercetin 3,7,3′-trimethyl ether, P3 6,8-Di-C-Methylquercetin 3,3′-dimethyl ether and the P4 6,8,3′-tri-C methylquercetin 3,7-dimethyl ether. Four known C-Methyl flavonols were also isolated from P. reticulatum for the first time and were P7 6-C-Methylquercetin 3-methyl ether, P8 6-C-Methylquercetin 3,7,3′-trimethyl ether, P9 6,8-Di-C-methylkaempferol 3-methyl ether and P10 6,8-Di-C-methylkaempferol 3,7-dimethyl ether. All of the isolated compounds were tested for their ability to inhibit prostaglandin synthesis and 6-CMethylquercetin 3,7,3′-trimethyl ether was found to be the most active (IC =6.58 mM) being about 300 times as 50 potent as aspirin.